A study showed that 64% of supermarket shoppers believe supermarket brands to be as good as national name brands. To investigate whether this result applies to its own product, the manufacturer of a national name-brand ketchup asked a sample of shoppers whether they believed that supermarket ketchup was as good as the national brand ketchup.

A.

If a sample of 100 shoppers showed 52 stating that the supermarket brand was as good as the national brand, what is the p-value?

Find the value of the test statistic. (Round your answer to two decimal places.)

Find the p-value. (Round your answer to four decimal places.)

p-value =
Answer #1

Solution :

This is the two tailed test .

The null and alternative hypothesis is

H0 : p = 0.64

Ha : p   0.64

= x / n = 52 / 100 = 0.52

Test statistic = z

= - P0 / [P0 * (1 - P0 ) / n]

= 0.52 - 0.64 / [(0.64 * 0.36) / 100]

= -2.5

P-value = 0.0124